OpenStudy (anonymous):

How much money should be invested at an annual rate of 12% in order to earn $300 interest in a year?

OpenStudy (ivettef365):

divide 300 / .12

OpenStudy (nurali):

I=Prt 300 = P * .12 * 1 P = 300/.12 $2500.00

OpenStudy (anonymous):

I= PRT/100 300= P x 12 x 1/100 P=2500

OpenStudy (anonymous):

thank yall